OpenStudy (anonymous):

Determine from your textbook or the Internet the equation for the rotational inertia I of a solid uniform disc of radius R, thickness h and mass M. You will use this expression to calculate the theoretical I for the disc used in this experiment. I was thinking I=mr^2?

OpenStudy (anonymous):

I = MR^2 should be right for a thin-walled hollow cylinder (think canned items). A hollow cylinder where the inner radius is small is I = 1/2 M(Inner Radius^2 + Outer Radius^2)

OpenStudy (anonymous):

what does is meant by solid uniform?

OpenStudy (anonymous):

@wastetime

OpenStudy (anonymous):

@Sukhvir Solid uniform means entirely solid, no holes or gaps which would make the disk a cylinder with no holes. Uniform states that the solid is the same density throughout the solid.

OpenStudy (anonymous):

@wastetime ok so why am i able to use I=mr^2 then?

OpenStudy (anonymous):

@Sukhvir I think I mistakenly wrote the wrong thing in my first post. I=MR^2 is the moment of inertia of a cylinder and a disk because a disk is a very thin cylinder.

OpenStudy (anonymous):

@wastetime Oh ok, so what is the formula of rotational interia that i would use for this type of question?

OpenStudy (anonymous):

I=mr^2/4?

OpenStudy (anonymous):

I=MR^2 should be right

OpenStudy (anonymous):

ok thanks @wastetime